Predict the direction in which the equilibrium wil
Answer #1

7. A.To the right as ka(HF) > Ka(C6H5COOH)

8. A. as only NH3 and Na2O yields basic solution when dissolved in water

9. pOH = -log[H+] = -log (6.7*10^-2) = 1.173

pH = 14- pOH = 12.83 (A)

10.6 S 2 1 ?1 ( EO AID (3.31 1-6)So,answer is C